OpenStudy (anonymous):

how to rewrite sqrt(2-sqrt(2)) * sqrt(2)?

OpenStudy (jdoe0001):

\(\bf \sqrt{2-\sqrt{2}} \times \sqrt{2}\) same roots multiply their content, so do that :)

OpenStudy (anonymous):

sqrt(4-sqrt(4))? :$

OpenStudy (jdoe0001):

\(\bf \sqrt{2-\sqrt{2}} \times \sqrt{2} \implies \sqrt{(2-\sqrt{2})\sqrt{2}}\)

OpenStudy (jdoe0001):

well, that's not ... hold on

OpenStudy (jdoe0001):

\(\bf \sqrt{2-\sqrt{2}} \times \sqrt{2} \implies \sqrt{(2-\sqrt{2})2}\)

OpenStudy (anonymous):

\[\sqrt{4-4\sqrt{2}} \] then? How would I simplify that further?

OpenStudy (jdoe0001):

well, 2 times 2 is 4, yes but \(\bf 2 \sqrt{2}\) is not \(\bf 4\sqrt{2}\)

OpenStudy (anonymous):

oh whoops! Yeah, my bad, lol.

OpenStudy (jdoe0001):

\(\bf 2 \times \sqrt{2}\) is just \(\bf 2\sqrt{2}\)

OpenStudy (jdoe0001):

and you can't simplify it further

OpenStudy (anonymous):

ah alright, thanks.

OpenStudy (jdoe0001):

yw